OpenStudy (anonymous):

Find the sum of the infinite geometric series, if it exists. 4 - 1 + 1/4 - 1/16 + . . .

OpenStudy (anonymous):

so r= -1/4 right

OpenStudy (anonymous):

\(a=4,r=-\frac{1}{4}\) use \[\frac{a}{1-r}\]

OpenStudy (anonymous):

yes, \(r=-\frac{1}{4}\)

OpenStudy (anonymous):

so the answer would be 16/5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\frac{4}{1+\frac{1}{4}}=\frac{4}{\frac{5}{4}}=\frac{16}{5}\] yes looks right